OpenStudy (anonymous):

What is the rate of change for f(x) = −2 cos 4x − 3 on the interval from x = pi over 4 to x = pi over 2?

OpenStudy (anonymous):

@aum negative pi over 16 negative 16 over pi 2 over pi pi over two

hartnn (hartnn):

use this formula \(\huge \dfrac{f(b)-f(a)}{b-a}\)

hartnn (hartnn):

basically you're finding the derivative, but over large interval, so we don't take limits. you know how to find f(pi/2) and f(pi/4) ?

OpenStudy (anonymous):

no nor really

OpenStudy (anonymous):

im not that good with trig

hartnn (hartnn):

to find f(pi/2), plug in x = pi/2 in f(x)

hartnn (hartnn):

to find f(pi/4), plug in x = pi/4 in f(x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so a is pi/2?

OpenStudy (anonymous):

and b is pi/4?

OpenStudy (anonymous):

or the other way around?

hartnn (hartnn):

doesn't matter :)

OpenStudy (anonymous):

ok so pi/8?

hartnn (hartnn):

pi/8 for what ?

OpenStudy (anonymous):

wait no pi/2! right?

OpenStudy (anonymous):

for the rate of change?

OpenStudy (anonymous):

bc before its simplified, its 4pi/8!!! and then simplify

OpenStudy (anonymous):

i foud the common denominator and ssimplifed

hartnn (hartnn):

so you got pi/4 in the denominator, right ?

hartnn (hartnn):

ok, f(pi/2) = -5 f(pi/4) = -1 numerator = -5-(-1) = -4 so, rate = -4/(pi/4) = -16/pi

OpenStudy (anonymous):

no i got 4pi/8

hartnn (hartnn):

thats incorrect how u got that ?

OpenStudy (anonymous):

common denominator by getting 8 and multiplying each by 4 and 2

OpenStudy (anonymous):

or would you hve to go to 16?

hartnn (hartnn):

8 so,